Information for Hobson Corp. for the current year ($ in millions):

Income from continuing operations before tax $ 155
Loss on discontinued operation (pretax) 32
Temporary differences (all related to operating income):
Accrued warranty expense in excess of expense
included in operating income
10
Depreciation deducted on tax return in excess of
depreciation expense
25
Permanent differences (all related to operating income):
Nondeductible portion of entertainment expense 5


The applicable enacted tax rate for all periods is 25%.

How much tax expense on income from continuing operations would be reported in Hobson's income statement?

  • $36.25 million.

  • $38.75 million.

  • $40.0 million.

  • $32.5 million.
Answer #1

Tax payable = (155 - 10 - 25 + 5) * 25% = 31.25 Million

Deferred tax = (10 + 25) * 25% = 35 * 25% = 8.75 million

Total tax expense = Tax payable + Deferred tax

= 31.25 million + 8.75 million = $40 million

Answer : $40.00 Million